### v1

# Nonlocal mass creation by metricized accelerations in Euclidean matterspace

Abstract
Relativistic four-acceleration in metrically organized motions can introduce the nonlocal mechanism for the local generation of active and reactive mass densities in the monistic geometrodynamics of correlated matterspace. Four Hilbert variations for geometrized matterspace with Euclidean subgeometry maintain the universal equivalence of active and reactive mass densities. The vector counterbalance in the contracted Bianchi identities with the acceleration-based curvatures prevents the gravitational collapse of non-local self-organizations.
Keywords Euclidean continuum, adaptive time, field monism, self-accelerating reality, geometrodynamics, non-local self-assembly
Mathematics Subject Classification 00A79, 83C99, 83E05
1 Introduction to field monism
In Europe, the idea of describing nature in terms of the properties of material space-plenum or ponderable ether goes back to the worldview of the ancient Greeks. “Plato in the Timaeus says that matter and space are the same” underlined Aristotle in his Physics [1]. Descartes’s matter-extension [2] from the seventeenth century was supported hundred years later by Lomonosov in his ether-based gravity due to local pushes by material liquid with invisible densities [3]. The pondarable space with inertial heat and ethereal masses was first used quantitatively by Umov in 1873 for his monistic cosmos of continuous energy flows, which can be reinterpreted by secondary relations with volumetric masses in the Newtonian empty space but not vice versa [4].
In Germany, Gustav Mie persistently supported the field theory of matter since his first publication on the continuous elementary particle [5]. It was in 1912, before the emerging era of quantum distributions in the non-local microworld. Einstein accepted the ponderable ether publicly in 1920 [6]. In 1929, the New Yorker Magazine popularized Einstein by noting “People slowly accustomed themselves to the idea that the physical state of space itself were finite physical reality”. For a long time, the monistic wave function for material densities of quantum particles was traditionally associated with the so-called microworld. The mechanical picture of the macroworld relied on Newtonian dualism for localized bodies and continuous forces (or massless fields) in empty space. This dualistic consept dominated at least until non-locality of superconducting currents was established in the laboratory and reliable measurements of megaparsec correlations in astronomy [7, 8].
In 1914, Einstein [9] proposed to nullify the geodesic four-acceleration $a_{\mu}\equiv u^{\nu}\nabla_{\nu}u_{\mu}\equiv u^{\nu}(\nabla_{\nu}u_{\mu}-\nabla_{\mu}u_{\nu})=u^{\nu}(\partial_{\nu}u_{\mu}-\partial_{\mu}u_{\nu})$ for the free fall of the probe body. By convention, the latter does not contribute to the external fields and the local derivatives $\nabla_{\nu}u_{\mu}\equiv\partial_{\nu}u_{\mu}-\Gamma_{\nu\mu}^{\rho}u_{\rho}$ with the symmetric affine correlations $\Gamma_{\nu\mu}^{\rho}=\Gamma_{\mu\nu}^{\rho}$. In the twenties century, the revolution 4-acceleration $a^{\nu}=g^{\mu\nu}a_{\mu}$ did not shake the particle-field dualism of matter. The point is that the relativistic mechanics and metric fields of Einstein, as well as the monistic energy transport of Umov, were satisfactorily referenced by the Newtonian formalism for many testable applications. But why does the dense mass of probe bodies not affect the external field generated universally by all masses? In this paper, we try to reconsider the relativistic four-vector $a^{\mu}$ as a primary notion in the mass-generation mechanism. But how to build geometrodynamics of inertial spaces with locally induced masses by metrically correlated accelerations in order to quantitatively describe nonlocal self-organization of mechanical macrodistributions with negligible dissipation losses?
In summer and autumn of 1915 Hilbert counseltad Einstein regarding the dual action with the classical Lagrangian for inertial sources and the Ricci scalar density for metric fields (around these sources). As a result, the Einstein Equation for ten tensor balances between metric fields and massive densities [10, 11] remained consistent with the Newtonian dualism for massive bodies and their forces or massless fields. In this dualistic approach to the monistic (quantum) reality, negative energies of gravitational or metric fields also remained as the independent notion, which ultimately led to the metric singularity and many mathematical surprises of post-Newtonian physics.
Despite the original intention of Hilbert to combine Mie’s theory of field matter [5] with Einstein’s relativistic ideas [9], the monistic macromechanics of inertial, massive fields was not the mainstream of theoretical studies during the last century. However, the 2022 Nobel Prize awarded in Physics for the nonlocal macro-experiments encouraged theorists to study long-range entanglement in the correlated macrocosm. Nonlocality in any closed hierarchical subsystem should survive until dissipative intervention or measurement, as in the mechanical microcosm of quantum nonlocality.
Nowadays the Cartesian mechanics of inertial space seems to have its renaissance due to new experimental data, including astro-correlations [7, 8]. This paper reveals covariant updates for the static equilibrium of the extended mass-energy, where the Ricci scalar density $R(r)\equiv(\mu_{a}+\mu_{p})8\pi G/c^{2}=4r_{o}^{2}/[r^{2}(r+r_{o})^{2}]$ is proportional to the equal densities of the active and passive (or reactive) masses [12]. The monistic macrophysics of inertial fields or field masses should clarify how inhomogeneous densities of kinetic energy resist gravitational collapse or resolve the Bentley paradox of Newtonian physics. The goal of our flatspace relativism with only four Hilbert variations under the six geometric bounds is to maintain Cartesian worldview quantitatively and to initiate the adaptive geometrodynamics of correlated inertial densities for further applications in nonlocal macrophysics, life sciences, engineering, and novel energo-informatics of big data.
2 Accelerations define field masses and monistic action
[…]
$\cases{u_{\mu}u^{\mu}\equiv 1,\ f^{\mu\nu}(x)\equiv g^{\mu\lambda}(x)g^{\nu\rho}(x)f_{\lambda\rho}(x)\equiv-f^{\nu\mu}(x),\cr f_{\mu\nu}(x)\equiv\nabla_{\mu}u_{\nu}(x)-\nabla_{\nu}u_{\mu}(x)=\partial_{\mu}u_{\nu}(x)-\partial_{\nu}u_{\mu}(x),\ \cr\nabla_{\mu}f_{\nu\lambda}(x)+\nabla_{\nu}f_{\lambda\mu}(x)+\nabla_{\lambda}f_{\mu\nu}(x)\equiv 0,\cr j^{\mu}(x)\!\equiv\!{c\varphi_{\!\!{}_{{}_{G}}}\partial_{\nu}\!\left[{\sqrt{-g(x)}}f^{\mu\nu}(x)\right]}/{4\pi\sqrt{-g(x)}},\ \nabla_{\mu}j^{\mu}(x)\equiv 0,\cr a^{\mu}(x)\equiv u_{\nu}f^{\nu\mu}(x),\ u^{\nu}\nabla_{\mu}u_{\nu}\equiv u_{\nu}\nabla_{\mu}u^{\nu}\equiv 0,\ u_{\mu}(x)a^{\mu}(x)\equiv\!0.}$ (1)
The static self-organization of the nonlocal charge in the laboratory takes the equilibrium energy density $\varphi_{\!{}_{G}}{\sqrt{G}}\mu(r^{\prime}=r)\equiv c^{2}\mu(r^{\prime}=r)$ around the point vertex at the coordinate origin. This static distribution follows the Poisson equation $\mu(r)=-\delta^{ij}\partial_{i}w_{j}/4\pi G$ with the metric field intensity $w_{i}(r)=\delta_{ij}w^{j}=-Gm(x^{j}/r)/[r(r+r_{o})]\equiv-c^{2}{\partial_{i}}ln{\sqrt{g_{oo}(r)}}$, $g_{oo}(r)=r^{2}/(r+r_{o})^{2}=1/g^{oo}(r)$ [12]. But inhomogeneous profiles of extended masses and their rest energies should be stabilized in the classical mechanics by some kind of non-Newtonian forces associated with the so-called Poincaré pressure [13]. How to balance the self-organizing forces in the monistic geometrodynamics of material space with affine connections or nonlocal correlations?
Again, the essential message of the radial solutions in [12] for static inertial distributions was the equivalence of the active (attractive) and passive (reactive) mass-energy densities, $\mu_{a}(r)c^{2}\equiv-c^{2}\partial_{i}w_{j}(r)\delta^{ij}/4\pi G$ $=w_{i}(r)w_{j}(r)\delta^{ij}/4\pi G\equiv\mu_{p}(r)c^{2}$. Both densities contribute to the Ricci scalar $R(r)=(\mu_{a}c^{2}+\mu_{p}c^{2})8\pi/\varphi_{\!{}_{G}}^{2}$ $=4r_{o}^{2}/[r(r+r_{o})]^{2}=2u_{o}i^{o}(r)-2f_{oi}(r)f^{oi}(r)$ in the co-moving frame with $u_{o}={\sqrt{g_{oo}(r)}}\equiv g_{o}(r)=r/(r+r_{o}),g_{i}\equiv g_{oi}/g_{o}=0,g_{oi}=0,g^{oi}=0,g^{ij}=-\delta^{ij}$, $g_{ij}=-\delta_{ij}$, $u_{i}=u^{i}=0$, $f_{oi}=-\partial_{i}g_{o}=-(x_{i}r_{o}/r)/(r+r_{o})^{2},f^{oi}=g^{oo}g^{ij}f_{oj}=\delta^{ij}x_{i}r_{o}/r^{3},(-f_{oi}f^{oi})=r^{2}_{o}/[r(r+r_{o})]^{2}=g_{o}i^{o},i^{o}=r_{o}^{2}/[r^{3}(r+r_{o})]$. The invariant sum of equal static contributions clarifies the tensor structure of the Lagrangian ${\cal L}(x)=-\varphi_{\!{}_{{}_{G}}}^{2}R(x)/{16\pi}$ or the scalar curvature $R(x)=2u_{\nu}(x)i^{\nu}(x)-f_{\mu\nu}(x)f^{\mu\nu}(x)>0$ in the monistic action $S_{m}$ for inertial (massive) fields by the local divergence of the coordinate self-accelerations $a^{\mu}(x)\equiv c^{2}u_{\nu}(x)f^{\nu\mu}(x)$:
$S_{m}\equiv-\!\int\!\frac{{\varphi_{\!{}_{{}_{G}}}^{2}}\!d^{4}x}{16\pi c}{\sqrt{-g}}\left(\!2u_{\nu}\!\nabla_{\mu}f^{\nu\mu}\!-\!f_{\mu\nu}\!f^{\mu\nu}\!\right)\!\equiv\!-\!\int\!\frac{{\varphi_{\!{}_{{}_{G}}}^{2}}d^{4}x}{8\pi c}\sqrt{-g}\nabla_{\mu}a^{\mu}$ (2)
$\equiv-\int\!\frac{{c^{3}}d^{4}x}{8\pi G}\sqrt{-g}[\nabla_{o}a^{o}(x)+\nabla_{i}a^{i}(x)]\equiv-\frac{c}{2}\int\!{{\sqrt{-g}}d^{4}x}[\mu_{p}(x)+\mu_{a}(x)].$ (3)
By defining the symmetric energy-momentum tensor $T_{\mu\nu}(x)$ from ${\sqrt{-g}}T_{\mu\nu}/2\equiv\partial({\sqrt{-g}}{\cal L})/\partial g^{\mu\nu}$ according to Hilbert [11] and using $\partial{\sqrt{-g}}\equiv-{\sqrt{-g}}g_{\mu\nu}\partial g^{\mu\nu}/2$ one gets the following densities of massive (inertial, material) fields:
$\cases{\frac{{\sqrt{-g}}}{2}T_{\mu\nu}\equiv\frac{c^{4}{\sqrt{-g}}}{16\pi G}(g_{\mu\mu}\nabla_{\lambda}a^{\lambda}-\nabla_{\mu}a_{\nu}-\nabla_{\nu}a_{\mu}),\cr T^{\nu}_{\mu}\equiv g^{\nu\lambda}T_{\mu\lambda}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}(\delta_{\mu}^{\nu}\frac{{B}}{2}-{B}_{\mu}^{\nu}),\ B_{\mu\nu}\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu},\cr{B}^{\nu}_{\mu}\equiv g^{\nu\lambda}B_{\mu\lambda},B^{\mu}_{\mu}\equiv 2\nabla_{\mu}a^{\mu}=R_{\mu}^{\mu}\Rightarrow\frac{4r_{o}^{2}}{r^{2}(r+r_{o})^{2}},a^{o}=0,u^{i}=0,\cr T^{\mu}_{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{\mu}a^{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{o}a^{o}\!+\!\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{i}a^{i}\equiv(\mu_{p}\!+\!\mu_{a})c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}(B_{o}^{o}\!+\!B^{i}_{i}),\cr\mu_{p}c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{o}a^{o}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}(\partial_{o}a^{o}+\Gamma_{o\lambda}^{o}a^{\lambda})\Rightarrow\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\Gamma_{oi}^{o}a^{i}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}r_{o}^{2}}{4\pi r^{2}(r+r_{o})^{2}},\cr\mu_{a}c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{i}a^{i}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}(\partial_{i}a^{i}+\Gamma_{i\lambda}^{i}a^{\lambda})\Rightarrow\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\partial_{i}a^{i}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}r_{o}^{2}}{4\pi r^{2}(r+r_{o})^{2}}.}$ (4)
The tensor curvature $B_{\mu\nu}\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu}$ $=R_{\mu\nu}+\theta_{\mu\nu}$, where $R_{\mu\nu}\equiv\partial_{\lambda}\Gamma^{\lambda}_{\mu\nu}-\partial_{\nu}\Gamma^{\lambda}_{\mu\lambda}+\Gamma^{\lambda}_{\mu\nu}\Gamma^{\rho}_{\rho\lambda}-\Gamma^{\lambda}_{\mu\rho}\Gamma^{\rho}_{\nu\lambda}$, $g^{\mu\nu}\theta_{\mu\nu}\equiv 0$, $g^{\mu\nu}B_{\mu\nu}=g^{\mu\nu}R_{\mu\nu}$ and $B_{\mu\nu}\neq R_{\mu\nu}$, is based in the novel geometric analyses on the covariant derivatives of four-acceleration $a_{\mu}$ under strict 3D-flattness, $g_{i}g_{j}-g_{ij}\equiv\delta_{ij},g_{\mu}\equiv g_{o\mu}/g_{o}$, $g^{ij}\equiv-\delta^{ij}$, and adaptive changes of local time, $d\tau\equiv g_{\mu}dx^{\mu}/c\neq dx^{o}/c$. For static self-organization with spherical symmetry, where $u_{i}\equiv\gamma(g_{i}-\beta_{i})=0,u^{i}\equiv\beta^{i}\gamma=0$, $g_{oi}=0$, $g^{oi}=0$, $\gamma\equiv 1/{\sqrt{1-\beta^{2}}}=1$, one can use the already derived [12] solution $g_{o}^{2}u^{o}=u_{o}\equiv g_{o}\gamma=g_{o}=r/(r+r_{o}),$ which results in $a^{o}\equiv u_{\nu}f^{\nu o}=0$, $a_{o}\equiv u^{\nu}f_{\nu o}=0$, $a^{i}\equiv u_{\nu}f^{\nu i}=u_{o}f^{oi}=u_{o}g^{o\lambda}g^{i\rho}f_{\lambda\rho}=u_{o}g^{oo}g^{ij}f_{oj}=\delta^{ij}\partial_{j}lng_{o}=\delta^{ij}\Gamma_{oj}^{o}=r_{o}x^{i}/[r^{2}(r+r_{o})]$ and $a_{i}\equiv u^{\nu}f_{\nu i}=u^{o}f_{oi}=-\partial_{i}lnu_{o}=-\Gamma^{o}_{oi}=-\delta_{ij}a^{j}$. Thus, the 2007 solution [12] for the strong-field interactions with probe (external) masses, $w_{i}(r)\equiv-Gmx_{i}/[r^{2}(r+r_{o})]\equiv a_{i}(r)c^{2}$, is determined by the Einsteinian 3-acceleration $a_{i}(r)$ of the static material densities within their correlated distribution in the volumetric mass integral.
The relativistic four-acceleration $a_{\mu}\equiv g_{\mu\nu}a^{\nu}$ is a primary characteristic of the local creation and dynamic self-distribution of massive densities in the nonlocal mass-energy $mc^{2}=const$. The action (2) with $\nabla_{\mu}a^{\mu}(x)$ would be dropped in the dualistic mechanics for localized masses near the coordinate origin due to the Gauss theorem for the distant hypersurface. In this traditional way, the mass-generation mechanism associated with $\nabla_{\mu}a^{\mu}(x)>0$ in the monistic mechanics of massive fields was dropped due to the assumed dualism of matter with massless geometric fields referenced by Newtonian empty space.
In the monistic mechanics for quasi-isolated (geometrized) distributions both geometric fields and inertial currents are defined by the local self-acceleration $a^{\mu}$ of matterspace with adaptive local times. The 3-acceleration $a^{i}\equiv g^{i\nu}a_{\nu}\equiv u_{\nu}f^{\nu i}\neq 0$ is responsible for everything in static self-organizations ($g^{oi}=0,g^{oo}=1/g_{oo}>1,a^{o}=0$, $a^{i}=-\delta^{ij}a_{j}$), including the generation of both active, $\mu_{a}=c^{2}\partial_{i}a^{i}/4\pi G$, and passive, $\mu_{p}=c^{2}a^{i}(\partial_{i}lng_{o})/4\pi G$, mass densities or Maxwell-type currents of mechanical energy. The spatial 3-divergence is well known as a local density of (active) charges in Coulomb forces.
In mechanics, the acceleration three-vector $a_{i}=-\delta_{ij}a^{j}$ works as the interaction field for the probe mass. The temporal divergence $\nabla_{o}a^{o}=\Gamma_{oi}^{o}a^{i}=-a_{i}a^{i}>0$ seems to be a local self-action associated with the metric time auto-changes $cd\tau/dx^{o}={\sqrt{g_{oo}}}=r/(r+r_{o})\neq const$. The Minkowski metric with $g_{oo}=1,g_{ij}=-\delta_{ij}$ has no local self-accelerations, $a_{o}=a^{o}=a_{i}=a^{i}=0$, $\nabla_{o}a^{o}=\nabla_{i}a^{i}=0$, and can produce neither the passive mass density $\mu_{p}$, nor the active $\mu_{a}$. Following the original initiative of Grossmann in 1912, the co-authors made only temporal changes in the Minkowski 4-interval, $ds=\sqrt{c^{2}d\tau^{2}(g_{\mu\nu})-\delta_{ij}dx^{i}dx^{j}}$, preserving Euclidean 3-space [14]. Metric time dilation can quantitatively explain all post-Newtonian corrections (light bending, Shapiro echo delay, perihelion precession, etc.) and to propose the acceleration mechanism for mass generation. It now appears that Grossmann reasonably resisted signing Einstein’s later manuscripts with the curvature of 3-space.
3 Non-local auto-dynamics of flat matterspace with adaptive time rate
The time-varying self-organization of correlated densities in nonlocal matterspace maintains everywhere $u_{\mu}u^{\mu}\equiv 1$ for the 4D geometry of 3+1 manifolds with adaptive time rates in the spatial continuum. Such self-consistent geometry never distorts the Euclidean 3D subgeometry in adaptive self-movements of the nonlocal unity. The theoretical requirement to preserve Euclidean 3-space by six inherent symmetries $g^{ij}\equiv-\delta^{ij}$ or $g_{ij}\equiv g_{i}g_{j}-\delta_{ij}$ reduces ten Hilbert variables $g^{\mu\nu}=g^{\nu\mu}$ [11] to four, $g^{\mu o}=g^{o\mu}$. One can eliminate the spatial variations in the Hilbert formalism, $T_{ij}\delta g^{ij}\equiv T_{ij}\delta(-\delta^{ij})=0$ in spite of $T_{ij}(x)\neq 0$.
Adaptive correlations in the geometrodynamics of continuous matterspace are maintained only by relevant responses of four metric potentials $g_{\mu}\equiv g_{oi}/g_{o}$ and the local time rate $d\tau\equiv g_{\mu}dx^{\mu}/c$, with $c^{2}d\tau^{2}\equiv g_{\mu\nu}dx^{\mu}dx^{\nu}+\delta_{ij}dx^{i}dx^{j}$, for all local changes within the constant integrals of volumetric mass-energy, 3-momentum, and angular momentum. There are no geometric restrictions for the time related variations, $\delta g^{\mu o}=\delta g^{o\mu}\neq 0$, in the flat-space formalism. Here one can expect only four independent equations after the metric variations, but not ten tensor balances as in the Einstein Equation for dualistic gravitation in the warped 3D void around Newtonian masses [10]. Four variational equations of continuous matterspace declare $T_{\mu o}=T_{o\mu}=0$ and leads quantitatively to the local equivalence of active and passive mass densities in all static and dynamic relations:
$\cases{T_{\mu o}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}(g_{\mu o}\nabla_{\lambda}a^{\lambda}-\nabla_{\mu}a_{o}-\nabla_{o}a_{\mu})\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{16\pi}(g_{\mu o}B-2B_{\mu o})=0,\cr g_{oo}{B}=2B_{oo}=4\partial_{o}a_{o}-4\Gamma_{oo}^{\lambda}a_{\lambda},\ g_{oi}{B}=2B_{oi}=2(\partial_{o}a_{i}+\partial_{i}a_{o})-4\Gamma_{oi}^{\lambda}a_{\lambda},\cr T^{\nu}_{o}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}(\delta^{\nu}_{o}\nabla_{\lambda}a^{\lambda}-\nabla_{o}a^{\nu}-\nabla^{\nu}a_{o})\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{16\pi}(\delta_{o}^{\nu}B-2B_{o}^{\nu})=0,\cr B_{o}^{i}\equiv g^{i\lambda}B_{o\lambda}\equiv\nabla_{o}a^{i}+g^{oi}\nabla_{o}a_{o}+g^{ji}\nabla_{j}a_{o}=0,\cr\nabla_{\lambda}a^{\lambda}\equiv\nabla_{o}a^{o}+\nabla_{i}a^{i}=2\nabla_{o}a^{o}=2\nabla_{o}a_{o}/g_{oo},\cr\nabla_{i}a^{i}\equiv B^{i}_{i}/2=B^{o}_{o}/2\equiv\nabla_{o}a^{o}\ $or$\ \mu_{a}(x)=\mu_{p}(x)\cr T^{o}_{o}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{16\pi}(B^{i}_{i}-B^{o}_{o})=0,\ B^{o}_{o}=B^{i}_{i}=\frac{1}{2}B\equiv\frac{1}{2}(B^{o}_{o}+B^{i}_{i})>0,\cr T^{i}_{i}=T^{\mu}_{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}B^{\mu}_{\mu}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}B^{i}_{i}\equiv 2\mu_{a}c^{2}=2\mu_{p}c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}B^{o}_{o}.}$ (5)
[…]
The massive field self-organization with $T^{\nu}_{o}(x^{i},x^{o})=T_{\nu}^{o}(x^{i},x^{o})=0$ and $T^{i}_{j}(x^{i},x^{o})\neq 0$ means that there is no 4D energy continuum for observations. Only 3D space $x^{i}$ is continuously filled by the measurable mass-energy, while $x^{o}$ is not the continuum coordinate. Rather, it is a world running parameter to form a specific metric time for each quasi-isolated system within the common spatial superposition under universal 3D geometry. If the external forces are neglected, the quasi-closed system is described by metric-kinetic flows of mass-energy in flat space. The corresponding components of the symmetric metric tensor $g_{\mu\nu}$ and the affine connections $\Gamma^{\lambda}_{\mu\nu}$ can be expressed in terms of the metric 4-potential $\varphi_{\!{}_{{}_{G}}}g_{\mu}$ with $g_{\mu}\equiv g_{o\mu}/{\sqrt{g_{oo}}}$:
$\cases{g_{ij}=g_{i}g_{j}-\delta_{ij},\ g^{ij}=-\delta^{ij},\ \delta^{ij}g_{i}=g_{o}g^{oj}\equiv g^{i},\ g^{oo}=(1-g_{i}g_{j}\delta^{ij})/g_{o}^{2},\cr\Gamma^{\lambda}_{\mu\nu}\equiv\frac{g^{\lambda\rho}}{2}(\partial_{\mu}g_{\nu\rho}+\partial_{\nu}g_{\mu\rho}-\partial_{\rho}g_{\mu\nu}),\cr\Gamma^{o}_{oo}=\partial_{o}lng_{o}+\delta^{ij}g_{i}(\partial_{o}g_{j}-\partial_{j}g_{o}),\ \Gamma^{i}_{oo}=\delta^{ij}g_{o}(\partial_{j}g_{o}-\partial_{o}g_{j}),\cr\Gamma^{o}_{oi}=(1-g_{k}g_{j}\delta^{kj})\partial_{i}lng_{o}-\frac{\delta^{kj}g_{k}}{2g_{o}}[\partial_{o}(g_{j}g_{i})+\partial_{i}(g_{o}g_{j})-\partial_{j}(g_{o}g_{i})],\cr\Gamma^{i}_{oj}=\delta^{ik}g_{k}\partial_{j}g_{o}-\frac{\delta^{ik}}{2}[\partial_{o}(g_{j}g_{k})+\partial_{j}(g_{o}g_{k})-\partial_{k}(g_{o}g_{j})],\cr\Gamma^{k}_{ij}\!=\!\frac{\delta^{kl}}{2}\!\left[\frac{g_{l}}{g_{o}}[\partial_{i}(g_{o}g_{j})\!+\!\partial_{j}(g_{o}g_{i})\!-\!\partial_{o}(g_{i}g_{j})]\!+\!\partial_{l}(g_{i}g_{j})\!-\!\partial_{i}(g_{l}g_{j})\!-\!\partial_{j}(g_{l}g_{i})\right],\cr\Gamma^{\nu}_{\mu\nu}=\frac{g^{\nu\rho}}{2}\partial_{\mu}g_{\nu\rho}=(1-g_{i}g^{i})\partial_{\mu}lng_{o}+\frac{g^{i}}{g_{o}}\partial_{\mu}(g_{o}g_{i})-\frac{\delta^{ij}}{2}\partial_{\mu}(g_{i}g_{j})=\partial_{\mu}lng_{o}.}$ (6)
Not arbitrary, but physically admissible transformations between 3+1 manifolds with inherent metric symmetries for Euclidean subgeometry, where $|g_{ij}(x)|$ $=1$ and ${\sqrt{{-g}(x)}}=g_{o}(x)$, are responsible for the tensor covariance in equations (4). The adaptive dilation $cd\tau(x)/dx^{o}\equiv g_{\mu}dx^{\mu}/dx^{o}$ of the physical time rate $d\tau$ in the geometric organization is not a primary entity to define the material processes, but a secondary function or observable phenomenon in the world picture and its description.
The covariant divergence of $T^{\nu}_{\mu}(x)$ should vanish in the Hilbert formalism due to the Noether theorem. The geometrized dynamics of the correlated continuum of inertial fields obeys the contracted Bianchi identity with the curvature tensor ${B}_{\mu\nu}(x)\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu}$ employed in terms of the relativistic four-accelerations $a_{\mu}(x)\equiv g_{\mu\nu}(x)a^{\nu}(x)$:
$\cases{\nabla_{\nu}T^{\nu}_{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}\nabla_{\nu}(\delta_{\mu}^{\nu}\frac{B}{2}-B_{\mu}^{\nu})\equiv\frac{c^{4}}{8\pi G}(\nabla_{\mu}\nabla_{\nu}a^{\nu}-\nabla_{\nu}\nabla_{\mu}a^{\nu}-\nabla_{\nu}\nabla^{\nu}a_{\mu})\cr\equiv-\frac{c^{4}}{8\pi G}(a^{\nu}R_{\nu\mu}+\nabla_{\nu}\nabla^{\nu}a_{\mu})=0,\cr a^{\nu}(\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu})+\nabla_{\nu}\nabla^{\nu}a_{\mu}=0,\ $if$\ a^{\nu}R_{\nu\mu}=a^{\nu}B_{\nu\mu},\ $and$\ \cr\nabla^{\nu}\nabla_{\nu}a_{o}=-2a^{o}\nabla_{o}a_{o},\ \nabla^{\nu}\nabla_{\nu}a_{i}=-a^{j}(\nabla_{i}a_{j}+\nabla_{j}a_{i})\ $for$\ B_{oj}=0.}$ (7)
One can rewrite the Euclidean space 3-balances in the 4-vector equation (6) in another equivalent form by using the geometric constraints ${\sqrt{-g(x)}}=g_{o}(x),g_{ij}(x)=g_{i}(x)g_{j}(x)-\delta_{ij}$ and the variational consequence $T^{\mu}_{o}(x)=T^{\mu o}(x)$ $=0$ for the correlated dynamics,
$\nabla_{\nu}T^{\nu}_{i}=\frac{\partial_{\nu}({\sqrt{-g}}T^{\nu}_{i})}{\sqrt{-g}}-\frac{T^{\rho\lambda}}{2}\partial_{i}g_{\rho\lambda}=\partial_{j}T^{j}_{i}+T^{j}_{i}\partial_{j}lng_{o}-\frac{T^{kj}}{2}\partial_{i}(g_{k}g_{j})=0.$ (8)
Inhomogeneous mass distributions cannot exist in static states in Newtonian mechanics and gravitation. These dualistic theories were developed for energy and momentum exchange between separated bodies in the void, not for nonlocal correlations of the elastic whole as a wave function continuum in quantum mechanics. The radial mass-energy densities of the static matterspace [12],
$\mu_{p}(r)c^{2}=\mu_{a}(r)c^{2}=\frac{c^{4}B(r)}{16\pi G}=\frac{Gm^{2}c^{4}}{4\pi r^{2}(c^{2}r+Gm)^{2}}\Rightarrow\frac{c^{4}}{4\pi Gr^{2}},$ (9)
has the universal curvature near the origin, $lim_{r\rightarrow 0}B(r\ll Gm/c^{2})$ $={4}/{r^{2}}$. For relatively large distances from the center of spherical symmetry, $r\gg Gm/c^{2}$, the centripetal gradients $c^{4}\partial_{i}B(r)/16\pi G<0$, are proportional to the mass integral $m$ and would predict the gravitational collapse for interacting partners or probe bodies.
In the monistic geometrodynamics of the nonlocal whole, there are no partners. Instead, there is a self-action of correlated densities. The corresponding self-accelerations, or counterforces $c^{4}\nabla_{\nu}B^{\nu}_{i}/8\pi G\equiv c^{4}(\nabla_{\nu}\nabla_{i}a^{\nu}+\nabla_{\nu}\nabla^{\nu}a_{i})/8\pi G$ after their substitution on the right hand side in (6), prevent the centripetal collapse of inhomogeneous energy profiles with $(\partial_{i}B)c^{4}/16\pi G\neq 0$. The covariant counterbalance for non-equilibrium oscillations around equilibrium profiles and for stationary organisations can be easily verified in (6)-(7) under the static self-organization of relativistic 4-velocities $cu_{\mu}\equiv g_{\mu\nu}cu^{\nu}\equiv\{cg_{o}\gamma;-c\gamma(\beta_{i}-g_{i})/g_{o}\}$ =$=\{c{\sqrt{g_{oo}}};0\}$ and 4-accelerations $a_{\mu}\equiv g_{\mu\nu}a^{\nu}\equiv c^{2}u^{\nu}f_{\nu\mu}=\{0;-c^{2}x_{i}r_{o}/[r^{2}(r+r_{o})]\}$, $a^{\mu}=\{0;+c^{2}x^{i}r_{o}/[r^{2}(r+r_{o})]\}$ of the generated mass-energy densities (8) with their metric consequences $g_{oo}=1/g_{oo}=r^{2}/(r+r_{o})^{2},g_{oi}=g^{oi}=0,g_{ij}=-\delta_{ij},g^{ij}=-\delta^{ij}$.
4 Shannon information in the yin-yang balance of distributed energies
The stabilizing balance of internal counter-forces in the extended charge was argued by Poincaré [13]. Our physical interpretation of the contracted Bianchi identities (5) through the dialectical unity and struggle of mechanical repulsions and attractions corresponds not only to Poincaré’s guidance, but also agrees with the stable cosmology of Kant. The latter was supported by the Hegelian ideas for the primacy of repulsion in the visible existence of volumetric matter (without the consequences of gravitational collapse or the Beantley paradox). Repulsion provides matter for attraction, as Hegel explained from the dialectical stability of material profiles on earth and in the cosmos. Below we mathematically decipher the Hehelean balance of repulsion and attraction in terms of positive (yang, active) and negative (yin, reactive) energies in the Shannon information continuum.
The Lagrange equation of motion (6) for the temporal axis, $\mu\Rightarrow 0$, in the flat-space physics (4) with $B^{i}_{o}=0$ and $B^{o}_{o}=B^{i}_{i}=B/2$,
$\cases{\nabla_{o}\frac{B}{2}\equiv\partial_{o}\frac{B}{2}=\nabla_{o}B^{o}_{o}+\nabla_{i}B^{i}_{o}\equiv[\partial_{o}B^{o}_{o}+\Gamma_{o\lambda}^{o}B^{\lambda}_{o}-\Gamma_{oo}^{\lambda}B_{\lambda}^{o}]\cr+[\partial_{i}B^{i}_{o}+\Gamma_{i\lambda}^{i}B^{\lambda}_{o}-\Gamma_{io}^{\lambda}B_{\lambda}^{i}]=\partial_{o}\frac{B}{2}+[\Gamma_{io}^{i}B^{o}_{o}-\Gamma_{io}^{j}B^{i}_{j}],\cr\Gamma_{io}^{j}[\nabla^{i}a_{j}+\nabla_{j}a^{i}]=\Gamma_{io}^{i}B^{o}_{o}=\Gamma_{io}^{i}B^{j}_{j}=2\Gamma_{io}^{i}\nabla_{j}a^{j},}$ (10)
can be read as the bounds for admissible 3-potentials $g_{i}\equiv g_{oi}/g_{o}$ in the relativistic condition $\Gamma_{io}^{j}B^{i}_{j}=\Gamma_{io}^{i}B/2$. Recall from (5) that $\Gamma_{io}^{j}=\Gamma_{io}^{i}=0$, if $g_{1}=g_{2}=g_{3}=0.$
Geometrodynamic equations (6)-(7) for three spatial axes, when $\mu\Rightarrow i$, can be simplified and solved analytically for the stationary self-organization of multi-vertex matterspace with $g_{o}(x)=g_{o}({\bm{x}}),g_{j}=0$, $a_{i}(\bm{x})=-\partial_{i}lng_{o}(\bm{x})=\partial_{i}ln[1/g_{o}(\bm{x})],a^{j}(\bm{x})=\delta^{jk}\partial_{k}lng_{o}(\bm{x})\equiv\partial^{j}lng_{o}(\bm{x})$:
$\cases{\nabla_{\nu}T^{\nu}_{i}=\frac{1}{g_{o}}{\partial_{j}(g_{o}T^{j}_{i})}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi c^{2}g_{o}}\{\partial_{i}[g_{o}(2\partial_{j}a^{j})]-\partial_{j}[g_{o}(\partial_{i}a^{j}-\delta^{jk}\partial_{k}a_{i})]\}=0,\cr\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}[(\partial_{i}lng_{o})\partial_{j}\partial^{j}lng_{o}+\partial_{i}\partial_{j}\partial^{j}lng_{o}]=\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}[(\partial_{j}lng_{o})\partial^{j}\partial_{i}lng_{o}+\partial_{j}\partial^{j}\partial_{i}lng_{o}],\cr\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\left(\partial_{i}ln\frac{1}{g_{o}(\bm{x})}\right)\partial_{j}\partial^{j}ln\frac{1}{g_{o}(\bm{x})}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\left(\partial_{j}ln\frac{1}{g_{o}(\bm{x})}\right)\partial^{j}\partial_{i}ln\frac{1}{g_{o}(\bm{x})},\cr\varphi_{\!{}_{{}_{G}}}^{2}{a_{i}}\partial_{j}a^{j}=\varphi_{\!{}_{{}_{G}}}^{2}{a_{j}}\partial^{j}a_{i},\ {\sqrt{G}}\mu_{pr}\varphi_{\!{}_{{}_{G}}}a_{i}\equiv f_{i}\equiv-{\sqrt{G}}\mu_{pr}\partial_{i}W,\ W\equiv-{\varphi_{\!{}_{{}_{G}}}}ln(\frac{1}{g_{o}}),\cr\frac{1}{g_{o}(\bm{x})}=1+\frac{r_{1}}{|{\bm{x}}-{\bm{\xi}}_{1}|}+\frac{r_{2}}{|{\bm{x}}-{\bm{\xi}}_{2}|}+...+\frac{r_{n}}{|{\bm{x}}-{\bm{\xi}}_{n}|},\ r_{k}\equiv{\sqrt{G}}m_{k}/\varphi_{\!{}_{{}_{G}}},\cr\varphi_{\!{}_{{}_{G}}}\{a_{1};a_{2};a_{3}\}=-\varphi_{\!{}_{{}_{G}}}\frac{\frac{r_{1}({\bm{x}}-{\bm{\xi}}_{1})}{|{\bm{x}}-{\bm{\xi}}_{1}|^{3}}+\frac{r_{2}({\bm{x}}-{\bm{\xi}}_{2})}{|{\bm{x}}-{\bm{\xi}}_{2}|^{3}}+...+\frac{r_{n}({\bm{x}}-{\bm{\xi}}_{n})}{|{\bm{x}}-{\bm{\xi}}_{n}|^{3}}}{1+\frac{r_{1}}{|{\bm{x}}-{\bm{\xi}}_{1}|}+\frac{r_{2}}{|{\bm{x}}-{\bm{\xi}}_{2}|}+...+\frac{r_{n}}{|{\bm{x}}-{\bm{\xi}}_{n}|}}\equiv-{\bm{\partial}}W(\bm{x}).}$ (11)
The metrically organized field intensity $\varphi_{\!{}_{{}_{G}}}\!{a_{i}}(\bm{x})$ for the probe charge ${\sqrt{G}}m_{pr}$ depends on the Shannon-type potential $W(\bm{x})=-\varphi_{\!{}_{{}_{G}}}ln[1+\varphi^{-1}_{\!{}_{{}_{G}}}\sum({\sqrt{G}}m_{k}/|{\bm{x}}-{\bm{\xi}}_{k}|)]$ for the optimal information transfer [16]. Shannon information exchange is embedded in the geometric organization of multi-vertex matterspace. Here the positive self-energy of active emergent charges $\varphi_{\!{}_{{}_{G}}}\sum{\sqrt{G}}m_{k}\equiv m_{k}c^{2}>0$ is compensated by the volumetric integral of the negative potential energies ${\sqrt{G}}\mu_{p}({\bf x})W({\bf x})<0$ of the reactive local densities ${\sqrt{G}}\mu_{p}({\bf x})$:
$\cases{{\sqrt{G}}\mu_{p}({\bf x})\equiv\frac{[-{\bm{\partial}}W({\bf x})]^{2}}{4\pi\varphi_{\!{}_{{}_{G}}}}=\frac{{\bm{\partial}}^{2}W({\bf x})}{4\pi}\equiv{{\sqrt{G}}\mu_{a}({\bf x})},\cr\int\!d^{3}x\!{\sqrt{G}}\mu_{a}({\bf x})\!\equiv\!\int\!\frac{\!d^{3}x\!}{4\pi}{\bm{\partial}}^{2}W({\bf x})\!=\!\!\int\!\frac{\!d{\bm{S}}\!}{4\pi}{{\bm{\partial}}W({\bf x})}\!=\!{\varphi}_{\!{}_{{}_{G}}}\!\sum_{k}\!r_{k}\!\equiv\!{\sqrt{G}}\sum_{k}\!m_{k},\cr\int\!\!{\sqrt{G}}\mu_{p}({\bf x})d^{3}x\!\equiv\!\frac{{\varphi}_{\!{}_{{}_{G}}}}{4\pi}\!\!\int\!\!\left(\!\frac{\frac{{({\bf x}-{\bf\xi}_{1})r_{1}}}{|{\bm{x}}-{\bm{\xi}}_{1}|^{3}}+\frac{({\bf x}-{\bm{\xi}}_{2})r_{2}}{|{\bf x}-{\bm{\xi}}_{2}|^{3}}+...+\frac{({\bf x}-{\bm{\xi}}_{n})r_{n}}{|{\bf x}-{\bm{\xi}}_{n}|^{3}}}{1+\frac{r_{1}}{|{\bf x}-{\bm{\xi}}_{1}|}+\frac{r_{2}}{|{\bf x}-{\bm{\xi}}_{2}|}+...+\frac{r_{n}}{|{\bf x}-{\bm{\xi}}_{n}|}}\!\right)^{\!2}\!\!\!d^{3}x=\!{\sqrt{G}}\sum_{k}\!m_{k},\cr\int\!\!{\sqrt{G}}\mu_{p}(\!{\bf x}\!)W\!(\!{\bf x}\!)d^{3}x\!\equiv\!-\!\!\int\!\!\frac{d^{3}x}{4\pi}\!\!\left(\!\!\frac{\varphi_{\!{}_{{}_{G}}}\!\!\sum_{k}\!\!\frac{{({\bf x}-{\bm{\xi}}_{k})r_{k}}}{|{\bf x}-{\bm{\xi}}_{k}|^{3}}}{1+\sum_{k}\!\!\frac{r_{k}}{|{\bf x}-{\bm{\xi}}_{k}|}}\!\!\right)^{2}\!\!\!\!ln\!\left(\!\!1\!+\!\sum_{k}\!\!\frac{r_{k}}{|{\bf x}-{\bm{\xi}}_{k}|}\!\!\right)\!=\!-\!c^{2}\!\sum_{k}\!\!m_{k},\cr\int\!d^{3}x[{\sqrt{G}}\mu_{a}({\bf x})\varphi_{\!{}_{{}_{G}}}+{\sqrt{G}}\mu_{p}({\bf x})W\!({\bf x})]=0.}$ (12)
The appearance of active charge densities ${\sqrt{G}}\mu_{a}(x)$ requires metric stresses and local energy costs for the generation of rest energy densities $\mu_{a}(x)c^{2}>0$ due to the positive self-potential $\varphi_{\!{}_{{}_{G}}}\equiv c^{2}/{\sqrt{G}}$. Positive rest energy of macroscopic bodies can be considered as kinetic and related to microscopic motions or fast micro-oscillations. Passive (or reactive) charge density ${\sqrt{G}}\mu_{p}(x)$ is affected by negative potential $W$ for information rates or metric compounds. The local energy imbalance $\mu_{a}(x)c^{2}+{\sqrt{G}}\mu_{p}(x)W(x)\neq 0$, where $\mu_{a}(x)=\mu_{p}(x)$, occurs under the volumetric zeroing in (11) of active (Yang, kinetic) and reactive (Yin, potential) self-energy integrals in the non-local distribution of matterspace with metric stresses and information exchanges.
In Europe, such yin-yang cancellation of kinetic and potential energies for a secluded star or other isolated system was first assumed by Pascual Jordan in 1936, but was not supported quantitatively by relevant calculations. Monistic geometrodynamics, where the flat matterspace of a quasi-isolated system relates the self-generated mass densities to local accelerations in the correlated distribution, returns to the integral balance of kinetic and potential energies under equal active and passive masses. The equivalence of passive (reactive) and active (attractive) mass densities is a mathematical consequence of four covariant relations $T_{o}^{\mu}(x)=0$ for the flat-space continuum of massive fields under the adaptive local time. The vector geometrodynamics of the correlated tensor densities $T_{i}^{j}(x)\neq 0$ in the nonlocal mass-energy organization obey the Bianci identities for relativistic four-accelerations. The latter are the primary reason for the observable existence of the physical world with inertial masses and adaptive changes of their densities.
5 Conclusion
Local dissipation processes from external (non-metric) interventions or intensive radiation losses of dense elastic regions violate the metric organization with nonlocal correlations in closed mechanical systems. Non-metric energy-momentum exchanges with $\nabla_{\mu}j^{\mu}(x)\neq 0$ are much easier to analyze in the dualistic terms of separated particles in Newton’s mechanical laws, rather than in the Cartesian terms of non-empty space with continuous vortex densities. The dualistic simplification of matter has had tremendous quantitative success in practice, leading to the Standard Model of physics for separated particles and their field mediators.
Despite the philosophical directives and the predicted evolution of physics towards the nondual field theory [15], one can say again that “people slowly accustomed themselves to the idea that the physical state of space itself were finite physical reality”. Cartesian publications on matter-extension and monistic physics are nevertheless still underrepresented in the current volume of dualistic alternatives for mechanical events.
The correlated matter-space or elastic space-plenum of the ancient Greeks can have many unique applications. It consistently accepts the nonlocal nature of elastic self-organization both in the quantum microcosm and in the whole Metagalaxy. Material space-plenum can be continuously filled with bits of information, which cannot be done for emptiness. Therefore, monistic physics has the power to initiate a new informatics of big data based on instantaneous stresses of the elastic cosmos, but not on the slow wave traffic of dissipative exchanges with retardation.
Practical applications of the monistic picture and the geometrodynamics of correlated distributions in macroscopic physics, biology, informatics, cosmology, hydraulic engineering and many other energy-related disciplines go beyond the presented extension of the static self-assembly [12] to its covariant self-motion in the absence of non-metric forces. Finally, the twentieth-century initiatives of Einstein-Grossmann, Mie, and Hilbert should one day return to the flat-space material plenum of the ancient Greeks and the monistic worldview of Descartes. The renaissance of Euclidean subgeometry for material space consistently explains the discovered entanglement of macroscopic states and can provide new testable predictions for further nonlocal experiments.
Data Availability Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest The author has no conflict of interest to declare that are relevant to the content of this article.